# Mapping the Optical Landscape of a Squaraine Molecule in the Visible and Ultraviolet Energy Range

### v1

@@ 2 Results and Discussion
In our investigation, we have specifically concentrated on the frontier orbitals, which show -character and are involved in the low-energy optical response. Analyzing our TDDFT and GW/BSE calculations, we have found a set of 12 orbitals, namely the 6 highest occupied and 6 lowest unoccupied molecular -orbitals, to be essential for capturing the excited states in the studied energy range. Accordingly, we have also chosen these orbitals (Fig. 2) as active space in the multi-configurational calculations. Besides the irreducible representation, we label the orbitals by their energetic ordering from DFT calculation counting the highest occupied (HOMO-n) and the lowest unoccupied molecular orbitals (LUMO+n). While the energetic order slightly varies with the level of theory, we observe no change in shape between the DFT and CASSCF-optimized orbitals. For more details, see Fig. S1.
@@ 2.1 Ground-State Absorption (GSA)
As a reference to previous investigations, we start by discussing data for the most prominent excitation of the spectrum at 1.92 eV in region I (Fig. 3(b)). Consistent with the 11, 41, we assign it to the first excited state S1, which has B3u symmetry (1-B3u), and predominantly arises from a HOMO (7b1u) LUMO (6b2g) transition with weights of 96% (TDDFT), 99% (GW/BSE), and 72% (CASSCF/NEVPT2). Contributions from all other transitions are below 5%.
When comparing the calculated with the experimental peak position at 1.92 eV, we observe an increasing improvement when climbing the theoretical ladder. While TDDFT (2.58 eV) clearly overestimates the experimental excitation energy (1.92 eV) by 0.61 eV, GW/BSE approaches the experimental value already closer (2.31 eV). Finally, CASSCF/NEVPT2 locates the peak at 2.10 eV.
@@
Moving on to region II, we address the four subtle signals around 3.5 eV (Fig. 3(c)). The interpretation of those proves to be intricate. Although our simulations locate in total four possible states in this energy range, only two of those are inherently symmetry-allowed. Firstly, there is the second state of symmetry, 2-, which mainly arises from a HOMO-4 (6) to LUMO (6) transition, admixed with minor double-excitations in the multi-configurational approach. Secondly, the neighboring state of symmetry, 2-, involves a transition between HOMO (7) and LUMO+1 (5). Their assignment to experimental features is complicated by an larger offset of the calculated excitation energies than encountered in region I. TDDFT and GW/BSE places the states at 4.44 and 4.47 eV (error of 0.8-1.2 eV) and 4.10 and 4.38 eV (error of 0.5-0.9 eV), respectively. With CASSCF, the states shift down to 4.00 and 4.39 eV. We estimate additional 0.09 eV from solvent effects, but still, the excitations are predicted notably too high in energy to unambiguously identify them as two of the experimental maxima (3.14, 3.45, 3.68 eV).
@@
As a second common prediction, all calculations show a significantly weaker signal 0.89 (TDDFT), 0.93 (GW/BSE) and 0.62 eV (CASSCF/NEVPT2) below the strong 6- absorption peak. It consists of two excitations, 3- and 4- placed closed to each other at about 5.67 (TDDFT), 5.47 (GW/BSE) and 5.25 eV (CASSCF/NEVPT2). Therefore, they are most likely responsible for the shoulder at 5.03 eV (E = 0.72 eV to experimental maximum) at the onset of region III. For all methods, one of the two states shows major contribution from a HOMO-7 (5) to LUMO (6) transition. Its counterpart is mainly described by an excitation from the HOMO-1 (3) to the LUMO+1 (5). Interestingly, their relative intensities depend on the level of theory, with TDDFT and GW/BSE favoring the first peak rather than the second as in CASSCF/NEVPT2.
@@ 2.2 Excited-State Absorption (ESA)
Concerning the measured spectrum, we want to emphasize that there are only two prominent features in the entire experimentally accessible window between 3 and 7 eV, namely at 4.38 eV and 4.78 eV. Their energy positions and relative intensities are in good agreement with previous reports46. In addition, our experimental data suggests a weak third contribution at 4.92 eV as a high-energy shoulder of the second peak. Note that the absence of absorption between 3.5 and 4 eV may be due to ground state depletion effects (1.9 eV ”pump” excitation to ” and 1.9 eV ”probe”), which precludes more definite statements about this energy range.
@@
To gain experimental insight into the orientation of the transition dipole moments for the two ESA transitions, and to deepen our understanding about the broadening of the second ESA peak with the potential shoulder at 4.92 eV, we determined the transient absorption anisotropy (TAA). We therefore measured the transient absorption for both parallel () and perpendicular () pump and probe polarizations and calculated the TAA value as discussed in the Methods section 4.1.3 50, 19. Note that erroneous values outside the range of -0.2–0.4 are possible in case of multiple contributing peaks of opposite sign 51, as well as sample degradation. This polarization-controlled TA measurement informs about the relative angle between the transition dipole moments of the pump and probe transitions, as detailed in the Methods Section.
@@
In contrast, for the region beyond 4.8 eV, data from both measurement approaches show a trend towards lower TAA. To explain this change, we have to include a third peak around 4.92 eV in our model, which is polarized orthogonal to the neighboring peak at 4.78 eV and has only one eighth of its intensity. We admit that these scans still suffer from relatively high uncertainties exceeding , reasons for which are discussed in TAA method section, owing to the noisy nature of the measured signals and possibly to some influence of a diffuse ground state depletion, which would be expected to increase with energy as can be seen in Fig. 3a, but did not result in observable peaks in this TA energy region.
We can think of two scenarios for the appearance of a high-energy shoulder in the second ESA peak. According to the first scenario, our TAA measurements reveal a change of polarization due to different electronic excitations. In this case, the rotation of the transition dipole from the long towards the short molecular axis would assign the shoulder to an excitation into a state. However, as a counterargument to this interpretation, we were not able to locate an electronic state with symmetry and comparable oscillator strength in this energy range in our calculations. In the second scenario, the shoulder in the ESA spectrum at 4.92 eV originates from a vibronic progression of its neighboring peak at 4.78 eV. This interpretation is supported by the energy difference of about 0.14 eV, which lies within the range expected for vibrational shoulders. According to the TA data, the vibration then would have to have a non-parallel transition dipole compared to the lower electronic 4- state.
@@ 2.3 Two-Photon Absorption (2PA)
Lastly, we want to complement our discussion with instantaneous two-photon absorption (2PA) data, offers additional evidence for the assignment of the electronic states. Recalling the selection rules of two-photon absorption, transitions from the fully-symmetric ground state Ag are allowed into other states with gerade parity, that is, Ag Ag and Ag Bg 52. Thus, in the case of our squaraine, the 2PA spectrum should coincidentally feature transitions into the same states as the excited state absorption process.
In Fig. 5, The two-photon absorption cross section of SQIB in toluene and acetonitrile is plotted as a function of the absolute photon energy (in eV). These data show that the choice of solvent does not noticeably affect the 2PA response of SQIB, as the peak positions remain essentially identical in both solvents. The measurement is based on two-photon fluorescence after excitation by an optical parametric amplifier. Although 2PA data for SQIB have been reported in the literature, the present measurements provide a more detailed view of the spectral response52. Two weak peaks at 2.06 and 2.27 eV are attributed to vibronic transitions of the first excited state . Although direct 2PA into this B3u state is symmetry-forbidden, weak vibronically allowed transitions can occur through coupling with odd-parity vibrational modes of Bu symmetry, which mix with the electronic B3u state and thus enable an effective Ag Ag transition 28, 52. The relative energetic distance of these features (0.21 eV) is reminiscent of the discussion of the ground-state absorption spectrum in Fig. S4(a). A more intense peak is observed around 3 eV, which we assign to the 2-Ag state. This assignment is consistent with the feature labeled as 2-Ag at 3.15 eV in Region II of the GSA spectrum discussed in Fig. 3(c). Finally, the highest measured point in the 2PA spectrum appears beyond 3.5 eV and has previously been estimated at 3.6 eV 11. Although the measurement could not be extended beyond this energy, it is reasonable to assume that the next transition would correspond to the 3-Ag state, and thus, we would expect the maximum to appear actually around 4.38 eV in accordance with our ESA analysis.
@@ 4.1.5 Two-Photon Absorption (2PA)
The two-photon absorption spectrum of SQIB in diluted toluene solution ( ) was collected adopting the two-photon excited fluorescence (2PEF) technique. The measurement was carried out using a Nikon A1R MP+ multiphoton upright microscope and a tunable (700-1300 nm) femtosecond mode-locked laser (Coherent Chameleon Discovery) as the excitation source. The excitation beam was focused on the sample, contained in a quartz fluorescence cuvette, through a 25x (NA = 1.1) water dipping objective. The outcoming 2PEF signal was collected in epifluorescence mode. The quadratic dependence of the two-photon absorption process on the excitation power was verified over all the excitation range.
@@ 5 Author contributions
MFS synthesized the SQIB compound and recorded the solution ground state spectra under the supervision of AL. MS prepared the PMMA-matrix samples and recorded the ground state transmission spectra and contributed to the manuscript writing. RS, FL, MJ and MK performed transient absorption measurements.
@@ 6 Note
The work is funded by the Austrian Science Fund via FWF Grants 10.55776/PAT1379523 and 10.55776/P36903. MS thanks the Linz Institute of Technology (LIT-2023-12-SEE-114 SLIME) for funding. MFS and AL are obliged to the DFG GRK 2591 “Template-designed Organic Electronics – TIDE” for financial support. MFS thanks the Manchot Foundation for a doctoral scholarship. AWH is grateful to NAWI Graz for support and acknowledges the use of HPC resources provided by the IT services of Graz University of Technology (ZID). We acknowledge ELI-ERIC for experimental and financial support during measurements at the ELI Beamlines under the projects ELIUPM4, TRANSDF and SQUAREPOP, especially Shirly Josefina Espinoza Herrera, Saul Vazquez Miranda and Mateusz Rebarz at the trELIps setup. We thank Anna Painelli for carefully reading the manuscript and providing helpful comments.

### v2

@@ 2 Results and Discussion
In our investigation, we have specifically concentrated on the frontier orbitals, which show -character and are involved in the low-energy optical response. Analyzing our TDDFT and GW/BSE calculations, we have found a set of 12 orbitals, namely the 6 highest occupied and 6 lowest unoccupied molecular -orbitals, to be essential for capturing the excited states in the studied energy range. Accordingly, we have also chosen these orbitals (Fig. 2) as an active space in the multi-configurational calculations. Besides the irreducible representation, we label the orbitals by their energetic ordering from DFT calculation counting the highest occupied (HOMO-n) and the lowest unoccupied molecular orbitals (LUMO+n). While the energetic order slightly varies with the level of theory, we observe no change in shape between the DFT and CASSCF-optimized orbitals. For more details, see Fig. S1.
@@ 2.1 Ground-State Absorption (GSA)
As a reference to previous investigations, we start by discussing data for the most prominent excitation of the spectrum at 1.92 eV in region I (Fig. 3(b)). Consistent with the references11, 41, we assign it to the first excited state S1, which has B3u symmetry (1-B3u), and predominantly arises from a HOMO (7b1u) LUMO (6b2g) transition with weights of 96% (TDDFT), 99% (GW/BSE), and 72% (CASSCF/NEVPT2). Contributions from all other transitions are below 5%.
When comparing the calculated with the experimental peak position at 1.92 eV, we observe an increasing improvement when climbing the theoretical ladder. While TDDFT (2.58 eV) clearly overestimates the experimental excitation energy (1.92 eV) by 0.61 eV, GW/BSE approaches the experimental value already closer (2.31 eV). Finally, CASSCF/NEVPT2 locates the peak at 2.20 eV.
@@
Moving on to region II, we address the four subtle signals around 3.5 eV (Fig. 3(c)). The interpretation of those proves to be intricate. Although our simulations locate in total four possible states in this energy range, only two of those are inherently symmetry-allowed. Firstly, there is the second state of symmetry, 2-, which mainly arises from a HOMO-4 (6) to LUMO (6) transition, admixed with minor double-excitations in the multi-configurational approach. Secondly, the neighboring state of symmetry, 2-, involves a transition between HOMO (7) and LUMO+1 (5). Their assignment to experimental features is complicated by an larger offset of the calculated excitation energies than encountered in region I. TDDFT and GW/BSE places the states at 4.44 and 4.47 eV (error of 0.8-1.2 eV) and 4.10 and 4.38 eV (error of 0.5-0.9 eV), respectively. With CASSCF/NEVPT2, the states shift down to 4.00 and 4.39 eV. We estimate additional 0.09 eV from solvent effects, but still, the excitations are predicted notably too high in energy to unambiguously identify them as two of the experimental maxima (3.14, 3.45, 3.68 eV).
@@
As a second common prediction, all calculations show a significantly weaker signal 0.89 (TDDFT), 0.93 (GW/BSE) and 0.62 eV (CASSCF/NEVPT2) below the strong 6- absorption peak. It consists of two excitations, 3- and 4- placed closed to each other at about 5.67 (TDDFT), 5.49 (GW/BSE) and 5.28 eV (CASSCF/NEVPT2). Therefore, they are most likely responsible for the shoulder at 5.03 eV (E = 0.72 eV to experimental maximum) at the onset of region III. For all methods, one of the two states shows major contribution from a HOMO-7 (5) to LUMO (6) transition. Its counterpart is mainly described by an excitation from the HOMO-1 (3) to the LUMO+1 (5). Interestingly, their relative intensities depend on the level of theory, with TDDFT and GW/BSE favoring the first peak rather than the second as in CASSCF/NEVPT2.
@@ 2.2 Excited-State Absorption (ESA)
Concerning the measured spectrum, we want to emphasize that there are only two prominent features in the entire experimentally accessible window between 3 and 7 eV, namely at 4.38 eV and 4.78 eV. Their energy positions and relative intensities are in good agreement with previous reports46. In addition, our experimental data suggests a weak third contribution at 4.93 eV as a high-energy shoulder of the second peak. Note that the absence of absorption between 3.5 and 4 eV may be due to ground state depletion effects (1.9 eV ”pump” excitation to ” and 1.9 eV ”probe”), which precludes more definite statements about this energy range.
@@
To gain experimental insight into the orientation of the transition dipole moments for the two ESA transitions, and to deepen our understanding about the broadening of the second ESA peak with the potential shoulder at 4.93 eV, we determined the transient absorption anisotropy (TAA). We therefore measured the transient absorption for both parallel () and perpendicular () pump and probe polarizations and calculated the TAA value as discussed in the Methods section 4.1.3 50, 19. Note that erroneous values outside the range of -0.2–0.4 are possible in case of multiple contributing peaks of opposite sign 51, as well as sample degradation. This polarization-controlled TA measurement informs about the relative angle between the transition dipole moments of the pump and probe transitions, as detailed in the Methods Section.
@@
In contrast, for the region beyond 4.8 eV, data from both measurement approaches show a trend towards lower TAA. To explain this change, we have to include a third peak around 4.93 eV in our model, which is polarized orthogonal to the neighboring peak at 4.78 eV and has only one eighth of its intensity. We admit that these scans still suffer from relatively high uncertainties exceeding , reasons for which are discussed in TAA method section, owing to the noisy nature of the measured signals and possibly to some influence of a diffuse ground state depletion, which would be expected to increase with energy as can be seen in Fig. 3a, but did not result in observable peaks in this TA energy region.
We can think of two scenarios for the appearance of a high-energy shoulder in the second ESA peak. According to the first scenario, our TAA measurements reveal a change of polarization due to different electronic excitations. In this case, the rotation of the transition dipole from the long towards the short molecular axis would assign the shoulder to an excitation into a state. However, as a counterargument to this interpretation, we were not able to locate an electronic state with symmetry and comparable oscillator strength in this energy range in our calculations. In the second scenario, the shoulder in the ESA spectrum at 4.93 eV originates from a vibronic progression of its neighboring peak at 4.78 eV. This interpretation is supported by the energy difference of about 0.14 eV, which lies within the range expected for vibrational shoulders. According to the TA data, the vibration then would have to have a non-parallel transition dipole compared to the lower electronic 4- state.
@@ 2.3 Two-Photon Absorption (2PA)
Lastly, we want to complement our discussion with two-photon absorption (2PA) data, offers additional evidence for the assignment of the electronic states. Recalling the selection rules of two-photon absorption, transitions from the fully-symmetric ground state Ag are allowed into other states with gerade parity, that is, Ag Ag and Ag Bg 52. Thus, in the case of our squaraine, the 2PA spectrum should coincidentally feature transitions into the same states as the excited state absorption process.
In Fig. 5, The two-photon absorption cross section of SQIB in toluene and acetonitrile is plotted as a function of the absolute photon energy (in eV). These data show that the choice of solvent does not noticeably affect the 2PA response of SQIB, as the peak positions remain essentially identical in both solvents. The measurement is based on two-photon excited fluorescence, i.e. the detection of light emitted after the simultaneous absorption of two photons. Although 2PA data for SQIB have been reported in the literature, the present measurements provide a more detailed view of the spectral response52. Two weak peaks at 2.06 and 2.27 eV are attributed to vibronic transitions of the first excited state . Although direct 2PA into this B3u state is symmetry-forbidden, weak vibronically allowed transitions can occur through coupling with odd-parity vibrational modes of Bu symmetry, which mix with the electronic B3u state and thus enable an effective Ag Ag transition 28, 52. The relative energetic distance of these features (0.21 eV) is reminiscent of the discussion of the ground-state absorption spectrum in Fig. S4(a). A more intense peak is observed around 3 eV, which we assign to the 2-Ag state. This assignment is consistent with the feature labeled as 2-Ag at 3.15 eV in Region II of the GSA spectrum discussed in Fig. 3(c). Finally, the highest measured point in the 2PA spectrum appears beyond 3.5 eV and has previously been estimated at 3.6 eV 11. Although the measurement could not be extended beyond this energy, it is reasonable to assume that the next transition would correspond to the 3-Ag state, and thus, we would expect the maximum to appear actually around 4.38 eV in accordance with our ESA analysis.
@@ 4.1.5 Two-Photon Absorption (2PA)
The two-photon absorption spectrum of SQIB in diluted toluene and acetonitrile solution ( ) was collected adopting the two-photon excited fluorescence (2PEF) technique. The measurement was carried out using a Nikon A1R MP+ multiphoton upright microscope and a tunable (700-1300 nm) femtosecond mode-locked laser (Coherent Chameleon Discovery) as the excitation source. The excitation beam was focused on the sample, contained in a quartz fluorescence cuvette, through a 25x (NA = 1.1) water dipping objective. The outcoming 2PEF signal was collected in epifluorescence mode. The quadratic dependence of the two-photon absorption process on the excitation power was verified over all the excitation range.
@@ 5 Author contributions
NT, PP, MK and AW conceived the research. NT and AW performed all calculations and analyzed them with contributons from AWH and PP. MFS synthesized the SQIB compound and recorded the solution ground state spectra under the supervision of AL. MS prepared the PMMA-matrix samples and recorded the ground state transmission spectra and contributed to the manuscript writing. RS, FL, MJ and MK performed transient absorption measurements. BB performed 2PA measurements. NT and AW wrote the manuscript with important contributions from all authors.
@@ 6 Note
The work is funded by the Austrian Science Fund via FWF Grants 10.55776/PAT1379523 and 10.55776/P36903. MS thanks the Linz Institute of Technology (LIT-2023-12-SEE-114 SLIME) for funding. MFS and AL are obliged to the DFG GRK 2591 “Template-designed Organic Electronics – TIDE” for financial support. MFS thanks the Manchot Foundation for a doctoral scholarship. AWH is grateful to NAWI Graz for support and acknowledges the use of HPC resources provided by the IT services of Graz University of Technology (ZID). BB acknowledges the equipment and framework of the COMP-R initiative, funded by the ”Departments of Excellence program of the Italian Ministry for University and Research” (MUR, 2023-2027). We acknowledge ELI-ERIC for experimental and financial support during measurements at the ELI Beamlines under the projects ELIUPM4, TRANSDF and SQUAREPOP, especially Shirly Josefina Espinoza Herrera, Saul Vazquez Miranda and Mateusz Rebarz at the trELIps setup. We thank Anna Painelli for carefully reading the manuscript and providing helpful comments.